# Limits of open ASEP stationary measures near a boundary

### v1

@@ 2. $\bullet$
holds at $(A,B,C,D)$ for some positive number $H=H(A,B,C,D)$ for $n>N$. Similar to the end of Step 2, for $n<N$, one can increase the value of $H$ if necessary so that $\mathsf{d}_{TV}\left(\mu_{n,m},\eta_{m}\right)$ can still be bounded by $\theta^{n}(Hm)^{3m}$ for any $1\leq m\leq n$. Step 3 is complete.
@@
In the rest of this subsection we will use Theorem 2.10 and the particle-hole duality to deduce Theorem 1.2 and Theorem 1.3. We recall the measures $\mu_{n|m}$ and $\mu_{n,m}$ from Remark 2.11. In view of the particle-hole duality stated as Lemma 2.7, we have:
@@
The next technical lemma shows that the total variation distance bound $\theta^{n}(Hm)^{3m}$ induces the correct growth rate of sequence $\{m_{n}\}_{n=1}^{\infty}$:
@@
$(Rm_{n})^{pm_{n}}\leq\left(\frac{Rsn}{\log n}\right)^{\frac{psn}{\log n}}=\left(\frac{Rs}{\log n}n\right)^{\frac{psn}{\log n}}\leq n^{\frac{psn}{\log n}}=n^{\frac{n\log\left(\frac{1}{\theta}\right)}{\log n}}=\left(n^{\frac{\log\left(\frac{1}{\theta}\right)}{\log n}}\right)^{n}=\left(n^{\log_{n}\left(\frac{1}{\theta}\right)}\right)^{n}=\left(\frac{1}{\theta}\right)^{n},$
@@
In the high density phase $A>1$, $A>C$, we assume $C/A\notin\{q^{l}:l\in\mathbb{Z}_{+}\}$ if $C\geq 1$ and analyze the measure $\eta_{m}$ defined in Theorem 2.10. In view of the second statement in Lemma 2.6, we have
@@
Before we provide the proof of Theorem 1.3, we present a special symmetry of multi-dimensional Askey–Wilson signed measures known as the `time reversal', which will be needed in the proof.
@@
In the low density phase $C>1$, $C>A$, we assume $A/C\notin\{q^{l}:l\in\mathbb{Z}_{+}\}$ if $A\geq 1$ and analyze the measure $\eta_{m}$ defined in Theorem 2.10. Notice $y_{0}(t)=\frac{1}{2}\left(\frac{C}{\sqrt{t}}+\frac{\sqrt{t}}{C}\right)$ and hence
@@
$\begin{split}&P_{1,t_{1}}^{(A,B)}\left(y_{0}(1),\mathsf{d}x_{1}\right)\dots P^{(A,B)}_{t_{m-1},t_{m}}\left(x_{m-1},\mathsf{d}x_{m}\right)\\ &=\pi_{t_{1}}^{(A,B,C,1/C)}(\mathsf{d}x_{1})P^{(A,B)}_{t_{1},t_{2}}\left(x_{1},\mathsf{d}x_{2}\right)\dots P^{(A,B)}_{t_{m-1},t_{m}}\left(x_{m-1},\mathsf{d}x_{m}\right)=\pi^{(A,B,C,1/C)}_{t_{1},\dots,t_{m}}\left(\mathsf{d}x_{1},\dots,\mathsf{d}x_{m}\right),\end{split}$
@@
where we have used (2.32) for $1\leq 1/t_{m}\leq\dots\leq 1/t_{1}<1+\varepsilon$ and the time reversal symmetry of Askey–Wilson signed measures (Proposition 2.19). The above identity coincides with (1.6) claimed in Theorem 1.3. By particle-hole duality, the total variation distance bound (1.7) in Theorem 1.3 directly follows from the bound (2.10) in Theorem 2.10. The last statement in the theorem (about the sequence $\{m_{n}\}$) follows directly from this bound combining with the technical Lemma 2.18. We conclude the proof. ∎
@@ Appendix A Bounding the total variation distance by generating functions
$\sum_{j\in\{0,1\}}\sum_{\upsilon_{1},\dots,\upsilon_{m-1}\in\{0,1\}}|G_{j}(t_{1,\upsilon_{1}},\dots t_{m-1,\upsilon_{m-1}})|\leq\frac{1+t_{m,1}}{t_{m,1}-t_{m,0}}\sum_{\upsilon_{1},\dots,\upsilon_{m}\in\{0,1\}}|G(t_{1,\upsilon_{1}},\dots t_{m,\upsilon_{m}})|.$ (A.3)
@@ Appendix B Total variation bounds of Askey–Wilson signed measures
Assume $A,C\geq 0$, $-1<B,D\leq 0$ and $q\in[0,1)$. Assume also $ABCD\notin\{q^{-l}:l\in\mathbb{N}_{0}\}$ and that $A/C\notin\{q^{l}:l\in\mathbb{Z}\}$ if $A,C\geq 1$. Then there exist positive constants $K\geq 1$ and $\varepsilon$ depending on $A,B,C$ and $q$ but not on $D$, such that for any $s<t$ in $I=[1,1+\varepsilon)$ and $x\in U_{s}$, the total variation of the Askey–Wilson signed measure $P_{s,t}(x,\mathsf{d}y)$ is bounded from above by $\frac{K}{(t-s)^{2}}$.
@@
As mentioned in Remark 2.13, the above total variation bound is more refined than the one utilized in [10]. In particular, the total variation bound of $P_{s,t}(x,\mathrm{d}y)$ provided by [10, Proposition A.1] can be derived as a simple corollary of the above result, but the converse does not hold. As we will see in the proof, more subtle estimates of the atom masses are needed to establish this result.
@@
Assume $q\in[0,1)$ and $(a,b,c,d)\in\Omega$, where $\Omega$ is defined in Definition 2.1. Then
@@
We first prove that there exists $K\geq 1$ and $\varepsilon>0$ depending on $A,B,C,D$ and $q$, such that for any $s<t$ in $I=[1,1+\varepsilon)$ and $x\in U_{s}$,
@@
At the end of the proof we will show that the above constant $K$ can be chosen as independent of $D$.
@@
The masses for atoms generated by $\mathbf{e}\in\{\mathbf{c},\mathbf{d}\}$ are given by similar formulas with $a$ and $\mathfrak{e}$ swapped. Note that $\mathbf{b}$ does not generate atoms since $b=B\sqrt{t}\in(-1,0]$. One can observe that the numerators of all atom masses are uniformly bounded from above. In the denominator, the term $(abcd)_{\infty}$ is uniformly bounded away from 00, since $abcd$ equals $ABt\leq 0$. Other three terms $(1-\mathfrak{e}^{2})$, $\mathfrak{e}^{4k}$ and $(q)_{k}$ are uniformly bounded away from 00, since we have $|\mathfrak{e}q^{k}|\geq 1$ and that $k\geq 1$ is bounded from above. Therefore, for any atom $y_{k}^{\mathbf{e}}$ generated by $\mathbf{e}\in\{\mathbf{a},\mathbf{c},\mathbf{d}\}$, we have:
@@
For fixed $\mathbf{e}$, $\mathbf{f}$ and $\kappa$, all the elements in the set $\mathcal{A}_{k}^{\mathbf{e},\mathbf{f}}$ are uniformly bounded away from 00 except for possibly one `exceptional' element. For this exceptional element $|1-q^{l}\mathfrak{f}/\mathfrak{e}|$, we have that either $\mathbf{f}=\mathbf{a}$ or $\mathbf{f}=\mathbf{c}$, and that it can be lower bounded by $(t-s)$ times a uniform positive constant.
If Claim B.4 holds, then for any fixed $\mathbf{e}\in\{\mathbf{a},\mathbf{c},\mathbf{d}\}$ and $k\in\mathbb{N}_{0}$, in view of (B.5), the atom mass $|p_{k}^{\mathbf{e}}(a,b,c,d)|$ is bounded by a uniform constant times $1/(t-s)^{2}$. Therefore in view of Lemma B.3, bound (B.2) holds. The proof of Claim B.4 necessitates a detailed analysis of all the possible values of $\mathfrak{f}/\mathfrak{e}$, which is provided in the following claim:
@@
Similar to the proof of Theorem 2.5, we consider the following three cases:
@@
    Let $x\in[-1,1]$ then $a=A\sqrt{t}$, $b=B\sqrt{t}$, $c=\sqrt{\frac{s}{t}}\left(x+\sqrt{x^{2}-1}\right)$ and $d=\sqrt{\frac{s}{t}}\left(x-\sqrt{x^{2}-1}\right)$. We have $|c|,|d|<1$ and $|c/a|=|d/a|=\sqrt{s}/(At)$.
@@
    Let $x=\frac{1}{2}\left(q^{j}A\sqrt{s}+\left(q^{j}A\sqrt{s}\right)^{-1}\right)$ for $j\in\mathbb{N}_{0}$ and $q^{j}A\sqrt{s}>1$. Then $a=A\sqrt{t}$, $b=B\sqrt{t}$, $c=q^{j}As/\sqrt{t}$ and $d=1/\left(q^{j}A\sqrt{t}\right)<1$. We have $c/a=q^{j}s/t$, $d/a=1/(q^{j}A^{2}t)$, $a/c=t/(q^{j}s)$ and $d/c=1/(q^{2j}A^{2}s)$. If $\mathfrak{e}=c$ then $q^{k}c=q^{k+j}As/\sqrt{t}\geq 1$, and hence $d/c\leq q^{2k}s/t$.
@@
    Let $x=\frac{1}{2}\left(q^{j}C/\sqrt{s}+\left(q^{j}C/\sqrt{s}\right)^{-1}\right)$ for $j\in\mathbb{N}_{0}$ and $q^{j}C/\sqrt{s}>1$. Then $a=A\sqrt{t}$, $b=B\sqrt{t}$, $c=q^{j}C/\sqrt{t}$ and $d=s/(q^{j}C\sqrt{t})<1$. We have $c/a=q^{j}C/(At)$, $d/a=s/(q^{j}ACt)$, $a/c=At/(q^{j}C)$ and $d/c=s/(q^{2j}C^{2})$. If $\mathfrak{e}=c$ then $q^{k}c=q^{k+j}C/\sqrt{t}\geq 1$, and hence $d/c\leq q^{2k}s/t$.
@@
    where $A\times B:=\{ab:a\in A,b\in B\}$ for $A,B\subset\mathbb{R}$. We factorize the numbers above as products of the parts that involve $A$ and $C$ and the other parts that involve $s$ and $t$ (which can be $1$):
    $\frac{\sqrt{s}}{At}=\frac{1}{A}\times\frac{\sqrt{s}}{t},\quad\frac{s}{t}=1\times\frac{s}{t},\quad\frac{t}{s}=1\times\frac{t}{s},\quad\frac{1}{A^{2}t}=\frac{1}{A^{2}}\times\frac{1}{t},\quad\frac{C}{At}=\frac{C}{A}\times\frac{1}{t},\quad\frac{At}{C}=\frac{A}{C}\times t.$
@@
    $\frac{\mathfrak{f}}{\mathfrak{e}}\in\left\{\frac{s}{t},\frac{t}{s}\right\}\times\{q^{r}:r\in\mathbb{Z}\}.$
    For $l\in\mathbb{Z}\setminus\{0\}$ we have $\left|1-q^{l}\frac{\mathfrak{f}}{\mathfrak{e}}\right|\geq\kappa$. For $l=0$ we have
    $\left|1-\frac{\mathfrak{f}}{\mathfrak{e}}\right|\geq\min\left(\left|1-\frac{s}{t}\right|,\left|1-\frac{t}{s}\right|\right)\geq\operatorname{Const}\times(t-s).$
@@
    $q^{-l}\in\left\{\frac{1}{A},\frac{1}{A^{2}},\frac{C}{A},\frac{A}{C}\right\}\quad\text{and}\quad\frac{\mathfrak{f}}{\mathfrak{e}}=q^{-l}\times\left\{\frac{\sqrt{s}}{t},\frac{1}{t},t\right\};$
    and for all other $l\in\mathbb{Z}$, $\left|1-q^{l}\frac{\mathfrak{f}}{\mathfrak{e}}\right|$ can be lower bounded by $\kappa$.

### v2

@@ 2. $\bullet$
holds at $(A,B,C,D)$ for some positive number $H=H(A,B,C,D)$ for $n\geq N$ and $1\leq m<n$. Similar to the end of Step 2, for $n<N$ or $m=n$, one can increase the value of $H$ if necessary so that $\mathsf{d}_{TV}\left(\mu_{n,m},\eta_{m}\right)\leq 1$ can still be bounded by $\theta^{n}(Hm)^{3m}$. Step 3 is complete.
@@
In the rest of this subsection we will use Theorem 2.11 and the particle-hole duality to deduce Theorem 1.2 and Theorem 1.4. We recall the marginal distributions $\mu_{n|m}$ and $\mu_{n,m}$ of $\mu_{n}$ and Remark 2.12. In view of the particle-hole duality stated as Lemma 2.7, we have:
@@
The next technical lemma shows that the total variation distance bound $\theta^{n}(Hm)^{3m}$ induces the correct growth rate of $\{m_{n}\}_{n=1}^{\infty}$. Note that we only need $p=3$ in this lemma.
@@
$(Rm_{n})^{pm_{n}}\leq\left(\frac{Rsn}{\log n}\right)^{\frac{psn}{\log n}}=\left(\frac{Rs}{\log n}n\right)^{\frac{psn}{\log n}}\leq n^{\frac{psn}{\log n}}\\ =n^{\frac{n\log\left(\frac{1}{\theta}\right)}{\log n}}=\left(n^{\frac{\log\left(\frac{1}{\theta}\right)}{\log n}}\right)^{n}=\left(n^{\log_{n}\left(\frac{1}{\theta}\right)}\right)^{n}=\left(\frac{1}{\theta}\right)^{n},$
@@
In the high density phase $A>1$, $A>C$, we assume $C/A\notin\{q^{l}:l\in\mathbb{Z}_{+}\}$ if $C\geq 1$ and analyze the measure $\eta_{m}$ introduced in Definition 2.10. In view of the second statement in Lemma 2.6, we have
@@
Before we provide the proof of Theorem 1.4, we present a special symmetry of multi-dimensional Askey–Wilson signed measures known as the ‘time reversal’, which will be needed in the proof.
@@
In the low density phase $C>1$, $C>A$, we assume $A/C\notin\{q^{l}:l\in\mathbb{Z}_{+}\}$ if $A\geq 1$ and analyze the measure $\eta_{m}$ introduced in Definition 2.10. Notice $y_{0}(t)=\frac{1}{2}\left(\frac{C}{\sqrt{t}}+\frac{\sqrt{t}}{C}\right)$ and hence
@@
$P_{1,t_{1}}^{(A,B)}\left(y_{0}(1),\mathsf{d}x_{1}\right)\dots P^{(A,B)}_{t_{m-1},t_{m}}\left(x_{m-1},\mathsf{d}x_{m}\right)\\ =\pi_{t_{1}}^{(A,B,C,1/C)}(\mathsf{d}x_{1})P^{(A,B)}_{t_{1},t_{2}}\left(x_{1},\mathsf{d}x_{2}\right)\dots P^{(A,B)}_{t_{m-1},t_{m}}\left(x_{m-1},\mathsf{d}x_{m}\right)\\ =\pi^{(A,B,C,1/C)}_{t_{1},\dots,t_{m}}\left(\mathsf{d}x_{1},\dots,\mathsf{d}x_{m}\right),$
@@
where we have used (2.32) for $1\leq 1/t_{m}\leq\dots\leq 1/t_{1}<1+\varepsilon$ and the time reversal symmetry of Askey–Wilson signed measures (Proposition 2.20). The above identity coincides with (1.6) in Definition 1.3. By particle-hole duality, the total variation distance bound (1.7) in Theorem 1.4 directly follows from the bound (2.10) in Theorem 2.11. The last statement in the theorem (about the sequence $\{m_{n}\}$) follows directly from this bound combining with the technical Lemma 2.19. We conclude the proof. ∎
2.3 Proof of Proposition 1.8
In this subsection we prove Proposition 1.8, which states that the limiting measure $\lambda$ on $\{0,1\}^{\mathbb{Z}_{+}}$ in the high density phase, as introduced in Definition 1.3, is a product Bernoulli measure if and only if $AC=1$, in which case it has density $A/(1+A)$.
Proof of Proposition 1.8.
Assume that the measure $\lambda$ on $\{0,1\}^{\mathbb{Z}_{+}}$ is product Bernoulli with density $\rho\in[0,1]$, then each $\lambda_{m}$ on $\{0,1\}^{m}$ for $m\in\mathbb{Z}_{+}$ is product Bernoulli with density $\rho$. Take $t_{1}=\dots=t_{m}=t\in(1-\varepsilon,1)$ in (1.6) we have
$\begin{split}\mathbb{E}_{\lambda_{m}}\left[t^{\tau_{1}+\dots+\tau_{m}}\right]&=(t\rho+1-\rho)^{m}\\ &=\frac{A^{m}}{(1+A)^{2m}}\int_{\mathbb{R}^{m}}\prod_{i=1}^{m}\left(1+t+2\sqrt{t}x_{i}\right)\pi^{(A,1/A,C,D)}_{t,\dots,t}\left(\mathsf{d}x_{1},\dots,\mathsf{d}x_{m}\right)\\ &=\frac{A^{m}}{(1+A)^{2m}}\int_{\mathbb{R}}\left(1+t+2\sqrt{t}x\right)^{m}\pi_{t}^{(A,1/A,C,D)}(\mathsf{d}x).\end{split}$ (2.33)
Recall that $\pi^{(A,1/A,C,D)}_{t}(\mathsf{d}x)=\nu\left(\mathsf{d}x;A\sqrt{t},\sqrt{t}/A,C/\sqrt{t},D/\sqrt{t}\right)$. Since $A>C\geq 0$, $A>1$ and $D\in(-1,0]$, one can shrink the value of $\varepsilon>0$ if necessary, so that for all $t\in(1-\varepsilon,1)$, we have $A\sqrt{t}>1$, $A\sqrt{t}>C/\sqrt{t}\geq 0$, $\sqrt{t}/A\in[0,1)$ and $D/\sqrt{t}\in(-1,0]$. Therefore the signed measure $\pi_{t}(\mathsf{d}x)$ only has atoms $>1$, and the largest atom is $y_{0}(t)=\frac{1}{2}\left(A\sqrt{t}+\frac{1}{A\sqrt{t}}\right)$. Using formula (2.3) one can check that $\pi_{t}(\{y_{0}(t)\})\neq 0$. We denote by $y_{1}^{*}(t)$ the second largest atom of $\pi_{t}(\mathsf{d}x)$ with nonzero mass, if it exists; otherwise, set $y_{1}^{*}(t)=1$. Then $y_{1}^{*}(t)<y_{0}(t)$, and that $\pi_{t}(\mathsf{d}x)$ is supported on $U_{t}\subset\{y_{0}(t)\}\cup[-1,y_{1}^{*}(t)]$. Therefore we have
$\int_{\mathbb{R}}\left(1+t+2\sqrt{t}x\right)^{m}\pi_{t}(\mathsf{d}x)=\\ \int_{\{y_{0}(t)\}}\left(1+t+2\sqrt{t}x\right)^{m}\pi_{t}(\mathsf{d}x)+\int_{-1}^{y_{1}^{*}(t)}\left(1+t+2\sqrt{t}x\right)^{m}\pi_{t}(\mathsf{d}x).$ (2.34)
The first term on the RHS equals $(1+t+2\sqrt{t}y_{0}(t))^{m}\pi_{t}(\{y_{0}(t)\})$. The absolute value of the second term on the RHS is bounded by $(1+t+2\sqrt{t}y_{1}^{*}(t))^{m}\left\lVert\pi_{t}\right\rVert_{TV}$, which has a lower order as $m\rightarrow\infty$ since $1<1+t+2\sqrt{t}y_{1}^{*}(t)<1+t+2\sqrt{t}y_{0}(t)$. Therefore by (2.33),
$\mathbb{E}_{\lambda_{m}}\left[t^{\tau_{1}+\dots+\tau_{m}}\right]=(t\rho+1-\rho)^{m}\sim\frac{A^{m}}{(1+A)^{2m}}\left(1+t+2\sqrt{t}y_{0}(t)\right)^{m}\pi_{t}(\{y_{0}(t)\}),$ (2.35)
where we write $f(m)\sim g(m)$ if $f(m)/g(m)\rightarrow 1$ as $m\rightarrow\infty$. The above implies
$t\rho+1-\rho=\frac{A}{(1+A)^{2}}\left(1+t+2\sqrt{t}y_{0}(t)\right)=\frac{1+At}{1+A}$
hence $\rho=A/(1+A)$, and also $\pi_{t}(\{y_{0}(t)\})=1$. Therefore, we conclude that (2.35) is an equality, and hence the second term on the RHS of (2.34) is equal to zero. If $y_{1}^{*}(t)>1$ is an atom, then by a similar analysis as above, this second term
$0=\int_{-1}^{y_{1}^{*}(t)}\left(1+t+2\sqrt{t}x\right)^{m}\pi_{t}(\mathsf{d}x)\sim\left(1+t+2\sqrt{t}y_{1}^{*}(t)\right)^{m}\pi_{t}(\{y_{1}^{*}(t)\}),$
which is a contradiction since $\pi_{t}(\{y_{1}^{*}(t)\})\neq 0$ by our assumption. Therefore we conclude that $\pi_{t}(\mathsf{d}x)$ only has a single atom at $y_{0}(t)$ with mass $1$. Observe that the continuous part density (2.2) of an Askey–Wilson signed measure has fixed sign over $x\in[-1,1]$, we conclude that the continuous part of $\pi_{t}(\mathsf{d}x)$ is constantly zero. Using formula (2.2) we have
$\left(q,t,AC,AD,C/A,D/A,CD/t\right)_{\infty}=0.$
Since $t\in(1-\varepsilon,1)$, $A>C\geq 0$ and $D\in(-1,0]$, we have $ACq^{k}=1$ for some $k\in\mathbb{N}_{0}$.
We first assume $k\in\mathbb{Z}_{+}$. We choose $t\in(1-\varepsilon,1)$ satisfying $A\sqrt{t}q\neq 1$. Then by $ACq^{k}=1$, we have either $A\sqrt{t}q>1$ or $C/\sqrt{t}>1$. When $A\sqrt{t}q>1$, by formula (2.4), the second largest atom generated by $A\sqrt{t}$, i.e., $\frac{1}{2}\left(A\sqrt{t}q+\frac{1}{A\sqrt{t}q}\right)$, has nonzero mass. When $C/\sqrt{t}>1$, by (2.3), the largest atom generated by $C/\sqrt{t}$, i.e., $\frac{1}{2}\left(\frac{C}{\sqrt{t}}+\frac{\sqrt{t}}{C}\right)$, has nonzero mass. This contradicts to the fact that $\pi_{t}(\mathsf{d}x)$ has only one atom $y_{0}(t)$ with nonzero mass.
We conclude that $k=0$, i.e., $AC=1$. When $AC=1$, by a computation, the multi-dimensional Askey–Wilson signed measure $\pi^{(A,1/A,C,D)}_{t_{1},\dots,t_{m}}\left(\mathsf{d}x_{1},\dots,\mathsf{d}x_{m}\right)$ on the RHS of (1.6) is a point mass at $x_{i}=y_{0}(t_{i})$ for $i=1,\dots,m$. Using (1.6) we conclude that $\lambda_{m}$ is the product Bernoulli measure on $\{0,1\}^{m}$ with density $A/(1+A)$. Therefore $\lambda$ is the product Bernoulli measure on $\{0,1\}^{\mathbb{Z}_{+}}$ with the same density. We conclude the proof. ∎
@@ Appendix A Bounding the total variation distance by generating functions
$\sum_{j\in\{0,1\}}\sum_{\upsilon_{1},\dots,\upsilon_{m-1}\in\{0,1\}}|G_{j}(t_{1,\upsilon_{1}},\dots t_{m-1,\upsilon_{m-1}})|\\ \leq\frac{1+t_{m,1}}{t_{m,1}-t_{m,0}}\sum_{\upsilon_{1},\dots,\upsilon_{m}\in\{0,1\}}|G(t_{1,\upsilon_{1}},\dots t_{m,\upsilon_{m}})|.$ (A.3)
@@ Appendix B Total variation bounds of Askey–Wilson signed measures
Assume $A,C\geq 0$, $-1<B,D\leq 0$ and $q\in[0,1)$. Assume also that $q^{l}ABCD\neq 1$ for all $l\in\mathbb{N}_{0}$ and that $A/C\notin\{q^{l}:l\in\mathbb{Z}\}$ if $A,C\geq 1$. Then there exist positive constants $K\geq 1$ and $\varepsilon$ depending on $A,B,C$ and $q$ but not on $D$, such that for any $s<t$ in $I=[1,1+\varepsilon)$ and $x\in U_{s}$, the total variation of the Askey–Wilson signed measure $P_{s,t}(x,\mathsf{d}y)$ is bounded from above by $\frac{K}{(t-s)^{2}}$.
@@
As mentioned in Remark 2.14, the above total variation bound is more refined than the one utilized in [30]. In particular, the total variation bound of $P_{s,t}(x,\mathrm{d}y)$ provided by [30, Proposition A.1] can be derived as a simple corollary of the above result, but the converse does not hold. As we will see in the proof, more subtle estimates of the atom masses are needed to establish this result.
@@
Assume $q\in[0,1)$ and $(a,b,c,d)\in\Omega$, where $\Omega$ is introduced in Definition 2.1. Then
@@
We assume that $A,C>0$ in the proof. Otherwise we are in the fan region $AC<1$, in which the Askey–Wilson signed measures are actually probability measures, whose total variation equal $1$. We first prove that there exists $K\geq 1$ and $\varepsilon>0$ depending on $A,B,C,D$ and $q$, such that for any $s<t$ in $I=[1,1+\varepsilon)$ and $x\in U_{s}$,
@@
At the end of the proof we will show that the above constants $K$ and $\varepsilon$ can be chosen as independent of $D$.
@@
The masses for atoms generated by $\mathbf{e}\in\{\mathbf{c},\mathbf{d}\}$ are given by similar formulas with $a$ and $\mathfrak{e}$ swapped. Note that $\mathbf{b}$ does not generate atoms since $b=B\sqrt{t}\in(-1,0]$. One can observe that the numerators of all atom masses are uniformly bounded from above. In the denominator, the term $(abcd)_{\infty}$ is uniformly bounded away from 00, since $abcd$ equals $ABt\leq 0$. Other three terms $(1-\mathfrak{e}^{2})$, $\mathfrak{e}^{4k}$ and $(q)_{k}$ are uniformly bounded away from 00, since we have $|\mathfrak{e}q^{k}|\geq 1$ and that $k\geq 1$ is bounded from above. Therefore, for any atom $y_{k}^{\mathbf{e}}$ generated by $\mathbf{e}\in\{\mathbf{a},\mathbf{c},\mathbf{d}\}$,
@@
For fixed $\mathbf{e}$, $\mathbf{f}$ and $k$, all the elements in the set $\mathcal{A}_{k}^{\mathbf{e},\mathbf{f}}$ are uniformly bounded away from 00 except for possibly one ‘exceptional’ element. For this exceptional element $|1-q^{l}\mathfrak{f}/\mathfrak{e}|$, we have that either $\mathbf{f}=\mathbf{a}$ or $\mathbf{f}=\mathbf{c}$, and that it can be lower bounded by $(t-s)$ times a uniform positive constant.
Notice that, although the RHS of (B.5) is an infinite product, we do not need to worry about its tail, because it is uniformly bounded as the inverse of a $q$-Pochhammer symbol. If Claim B.4 holds, then for any fixed $\mathbf{e}\in\{\mathbf{a},\mathbf{c},\mathbf{d}\}$ and $k\in\mathbb{N}_{0}$, there are at most two pairs of $\mathbf{f}\in\{\mathbf{a},\mathbf{c}\}$ and $l\geq-k$ that produce exceptional elements $|1-q^{l}\mathfrak{f}/\mathfrak{e}|$. Hence in view of (B.5), the atom mass $|p_{k}^{\mathbf{e}}(a,b,c,d)|$ is bounded by a uniform constant times $1/(t-s)^{2}$. Therefore in view of Lemma B.3, bound (B.2) holds.
The proof of Claim B.4 necessitates a detailed analysis of all the possible values of $\mathfrak{f}/\mathfrak{e}$, which is provided in the following claim:
@@
We recall that we always have $a>0$ and $b\in(-1,0]$. Similar to the proof of Theorem 2.5, we consider the following three cases depending on $x\in U_{s}$:
@@
    Let $x\in[-1,1]$ then we have $a=A\sqrt{t}$, $b=B\sqrt{t}$, $c=\sqrt{\frac{s}{t}}\left(x+\sqrt{x^{2}-1}\right)$ and $d=\sqrt{\frac{s}{t}}\left(x-\sqrt{x^{2}-1}\right)$. We have $|c|,|d|<1$ and $|c/a|=|d/a|=\sqrt{s}/(At)$. Claim B.5 holds.
@@
    Let $x=\frac{1}{2}\left(q^{j}A\sqrt{s}+\left(q^{j}A\sqrt{s}\right)^{-1}\right)$ for $j\in\mathbb{N}_{0}$ and $q^{j}A\sqrt{s}>1$. Then $a=A\sqrt{t}$, $b=B\sqrt{t}$, $c=q^{j}As/\sqrt{t}>0$ and $d=1/\left(q^{j}A\sqrt{t}\right)\in(0,1)$. We have $c/a=q^{j}s/t$, $d/a=1/(q^{j}A^{2}t)$, $a/c=t/(q^{j}s)$ and $d/c=1/(q^{2j}A^{2}s)$. If $\mathfrak{e}=c$ then $q^{k}c=q^{k+j}As/\sqrt{t}\geq 1$, and hence $d/c\leq q^{2k}s/t$. Therefore Claim B.5 holds.
@@
    Let $x=\frac{1}{2}\left(q^{j}C/\sqrt{s}+\left(q^{j}C/\sqrt{s}\right)^{-1}\right)$ for $j\in\mathbb{N}_{0}$ and $q^{j}C/\sqrt{s}>1$. Then $a=A\sqrt{t}$, $b=B\sqrt{t}$, $c=q^{j}C/\sqrt{t}>0$ and $d=s/(q^{j}C\sqrt{t})\in(0,1)$. We have $c/a=q^{j}C/(At)$, $d/a=s/(q^{j}ACt)$, $a/c=At/(q^{j}C)$ and $d/c=s/(q^{2j}C^{2})$. If $\mathfrak{e}=c$ then $q^{k}c=q^{k+j}C/\sqrt{t}\geq 1$, and hence $d/c\leq q^{2k}s/t$. Therefore Claim B.5 holds.
@@
    where we denote $A\times B:=\{ab:a\in A,b\in B\}$ for $A,B\subset\mathbb{R}$. We factorize the numbers above as products of the parts that involve $A$ and $C$ and the other parts that involve $s$ and $t$ (which can be $1$):
    $\frac{\sqrt{s}}{At}=\frac{1}{A}\times\frac{\sqrt{s}}{t},\quad\frac{s}{t}=1\times\frac{s}{t},\quad\frac{t}{s}=1\times\frac{t}{s},\quad\\ \frac{1}{A^{2}t}=\frac{1}{A^{2}}\times\frac{1}{t},\quad\frac{C}{At}=\frac{C}{A}\times\frac{1}{t},\quad\frac{At}{C}=\frac{A}{C}\times t.$
@@
    $\frac{\mathfrak{f}}{\mathfrak{e}}\in\left\{\frac{s}{t},\frac{t}{s}\right\}\times\{q^{r}:r\in\mathbb{Z}\}:$
    Denote $\frac{\mathfrak{f}}{\mathfrak{e}}\in\left\{q^{h}\frac{s}{t},q^{h}\frac{t}{s}\right\}$, $h\in\mathbb{Z}$. Then if $h+l\in\mathbb{Z}\setminus\{0\}$ we have $\left|1-q^{l}\frac{\mathfrak{f}}{\mathfrak{e}}\right|\geq\kappa$. If $h+l=0$,
    $\left|1-q^{l}\frac{\mathfrak{f}}{\mathfrak{e}}\right|\geq\min\left(\left|1-\frac{s}{t}\right|,\left|1-\frac{t}{s}\right|\right)\geq\operatorname{Const}\times(t-s).$
@@
    $q^{-l}\in\left\{\frac{1}{A},\frac{1}{A^{2}},\frac{C}{A},\frac{A}{C}\right\}\quad\text{and}\quad\frac{\mathfrak{f}}{\mathfrak{e}}=\left\{\frac{\sqrt{s}}{t},\frac{1}{t},t\right\}\times q^{-l};$
    and for all other $l\in\mathbb{Z}$, $\left|1-q^{l}\frac{\mathfrak{f}}{\mathfrak{e}}\right|$ can be lower bounded by $\kappa$. The above fact follows from
    $\min\left(\left|1-\frac{\sqrt{s}}{t}\right|,\left|1-\frac{1}{t}\right|,|1-t|\right)\geq\operatorname{Const}\times(t-s),$
    and from our choices of $\varepsilon>0$ and $\kappa>0$.